OpenStudy (anonymous):

the height of one square pyramid is 6 m. A similar pyramid has a height of 12 m. The volume of the smaller pyramid is 400 m^3. A. Scale Factor of the smaller pyramid to the larger pyramid in simplest form B. Ratio of the areas of the bases of the smaller pyramid to the larger pyramid. C. Ratio of the volume of the smaller pyramid to the larger pyramid. D. Volume of the larger pyramid.

OpenStudy (sidsiddhartha):

scale factor=height of pyramid A/height of pyramid B=12/6=2:1 now area=2^2:1^2=4:1 volume=2^3:1^3=8:1 now u know the volume ratio so use that

OpenStudy (anonymous):

Then the volume of the larger pyramid is 400 x 8 which is 3200m^3 right? :)

OpenStudy (anonymous):

Isn't that only if the larger pyramid's volume was 400 m^3? This question says that the smaller pyramid is 400 m^3...Sorry I'm just really lost.

OpenStudy (sidsiddhartha):

sorry yeah ure right i misread it :)

OpenStudy (anonymous):

So it is 400 x 8 right? :D

OpenStudy (sidsiddhartha):

yeah :)

OpenStudy (anonymous):

Okay thank you so much ^^ <3

OpenStudy (sidsiddhartha):

^_^